# Dynamic Mixture of Progressive Parameter-Efficient Expert Library for Lifelong Robot Learning

### v1

@@ A.1 Policy and Algorithm Details
As illustrated in Section 4.1, we use an auto-regressive policy consisting of vision, text, and state encoders, an input modality fusion module, a temporal transformer, and an action head [40, 43]. Detailed hyperparameters of each submodule are summarized in Table 1.
We utilize the implementations of Experience Replay (ER) [22], Elastic Weight Consolidation (EWC) [27], and PackNet [30] as provided in LIBERO [43]. For ER, we employ a replay ratio of 20%. For EWC, we set the hyperparameters $\gamma=0.9$ and $\lambda=5\times 10^{4}$. For PackNet, we fix 25% of the parameters as important. When training with ER and EWC, we use two GPUs, each with a batch size of 16, ensuring the total batch size remains consistent with other methods. For additional details on these baselines, we refer readers to [43]. We directly use the official code of LOTUS [92]. Since LOTUS employs a distinct architecture, we adjust its parameters to be comparable to our policy and follow the same training pipeline. Additionally, LOTUS utilizes the DINO-v2 [14] foundation model instead of CLIP, as recommended by the ablation study in their paper.
For LoRA-based methods, we insert LoRA modules with rank $r=8$ into the 6-th to 11-th layers of the CLIP image and text encoders, and LoRA modules with $r=16$ into the temporal transformer. LoRA is applied to the query and value projection matrices. In TAIL and IsCiL, we learn task-specific fusion modules, state encoders, and action heads, as described in [40]. In contrast, these components are frozen in L2M following the original paper [41]. Due to the unavailability of publicly released code for TAIL, we implemented it based on the description in the paper [40]. We integrated the code for L2M and IsCiL into our codebase, as the original implementations operate with different policies across different benchmarks. For L2M, we use a pool size of 30 and the similarity loss coefficient $\lambda=0.5$. For IsCiL, we set the skill prototype $\mathcal{X}_{z}$ in $\mathcal{X}$ to be composed of 20 bases as in [42]. For DMPEL, we adopt the same LoRA settings in the image encoder, text encoder, and temporal transformer, and use LoRA experts with $r=16$ for all linear layers in other submodules. Besides, in the Long suite, we tune both the low-rank weights $\boldsymbol{A},\boldsymbol{B}$ and biases $\boldsymbol{b}$, whereas in the Goal, Spatial, and Object suites, we only tune the low-rank weights $\boldsymbol{A},\boldsymbol{B}$. However, both settings lead to fewer learnable parameters than TAIL [40], which learns these sub-modules separately for each task. We also employ dropout with probability $p=0.15$ on the coefficient for low-rank experts to avoid overfitting.
Table 1: Environment and Policy Architecture Configuration
Hyperparameter Value Hyperparameter Value
@@
Pretrained model CLIP ViT-B/16 [15] Image resolution 128*128
@@
Joint state dim. 7 Gripper state dim. 2
@@
Number of hidden layers 1 Hidden size 256
@@
Number of layers 6 Number of attn. heads 8
Embedding size 768 MLP hidden size 1024
Max sequence Length 10 Dropout 0.15
@@
Number of hidden layers 2 Hidden size 256
Number of Modes 5 Min. std 1e-4
Activation Softplus Action dim. 7
@@
Number of hidden layers 2 Hidden Activation GELU
@@ A.2 Training Configuration
We generally follow the setup in the LIBERO benchmark for training and evaluation [43]. Detailed settings and hyperparameters are summarized in Table 2. For all experiments, we use either Nvidia A100 or H800 GPUs. We employ Distributed Data Parallel (DDP) for parallel training across 4 GPUs during pretraining, while using a single GPU for lifelong learning. We utilize the 16-bit floating point precision (FP16) training mode to accelerate the training process. To ensure reproducibility, we adopted three random seeds, 100, 200, 300, and we report the mean and standard deviation of the performance over three seeds throughout the paper.
Table 2: Training Hyperparameters
Hyperparameter Value Hyperparameter Value Hyperparameter Value
Optimizer
Training epochs 10 Batch size 32 Optimizer AdamW
Betas (0.9,0.999) Learning rate 1e-4 Anneal strategy cos
Weight decay 0.1 Gradient clip 100
Image Augmentation
Brightness 0.3 Saturation 0.3 Contrast 0.3
Hue 0.3 Color jitter prob. 0.9 Rotation 15
Translation 0.1 Affine prob. 0.6
Evaluation
Epochs per eval. 2 Eval. episodes 20
A.3 Benchmark Details
We present the language instructions for the tasks in four lifelong learning LIBERO task suites in Table 3. Although some tasks have similar descriptions, they are not identical due to differences in environment configurations, such as varying spatial layouts, objects, or goal positions. Each task provides 50 successful expert demonstrations for imitation learning.
Table 3: Taks Instructions of LIBERO Task Suites
Suite Instructions
@@ A.3 Benchmark Details
Goal put the bowl on top of the cabinet
@@
Spatial pick up the black bowl in the top drawer of the wooden cabinet and place it on the plate
@@
Object pick up the ketchup and place it in the basket
@@
Long put the white mug on the left plate and put the yellow and white mug on the right plate
@@ Appendix B Additional Results
B.1 Learning curves
We present the learning curves for all methods in four lifelong learning LIBERO task suites in Figure 7-10 . The $y$-axis indicates the success rate, while the $x$-axis depicts the agent’s training process over the course of lifelong learning. For example, the Task 1 subplot in the figure shows the agent’s performance on the first task as it learns ten tasks sequentially.

### v2

@@ A.1 Policy and Algorithm Details
We use a policy consisting of vision, text, and state encoders, an input modality fusion module, a temporal transformer, and an action head (Liu et al., 2024b; a). We utilize the open-source CLIP ViT/B-16 model (Radford et al., 2021) as vision and text encoders, while all other components of the policy is learned from robot demonstrations. Detailed hyperparameters of each sub-module are summarized in Table 2, most of them are naturally inherited from the open-source LIBERO benchmark and CLIP model.
Baselines.
We consider naive sequential fine-tuning baselines, representative lifelong learning methods that involve tuning the entire policy, and several LoRA-based methods:
  • Sequential Fine-tuning (SeqFT) (Liu et al., 2024a): Naively perform full fine-tuning (FFT) or use frozen pretrained feature (FPF), i.e., freeze the spatial encoders and fine-tune the rest, on sequentially arriving tasks.
  • Experience Replay (ER) (Chaudhry et al., 2019), a replay method that maintains a memory buffer with samples from previous tasks. We store a subset of data after each task (20% in our experiments) and uniformly sample a fixed number of replay data from the buffer along with new task data during training. We use a batch size of 32 for old data, consistent with the batch size for new data, and mix them up for every training iteration.
  • Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017), a regularization method that uses the Fisher Information Matrix (FIM) to quantify the importance of each parameter and adds a regularization term to the learning objective to constrain network updates. We set the hyperparameters $\gamma=0.9$ and $\lambda=5\times 10^{4}$.
  • PackNet (Mallya & Lazebnik, 2018), an architectural method that first trains and prunes the network to retain a fixed proportion (25% in our experiments) of the most important parameters. The selected parameters are then fine-tuned and frozen to prevent forgetting.
  • LifelOng knowledge Transfer Using Skills (LOTUS) (Wan et al., 2024) involves two stages: extracting temporal segment features by leveraging frozen foundation models for skill discovery, and hierarchical imitation learning with the meta-controller and a growing skill library. Since LOTUS employs a distinct architecture, we adjust its parameters to be comparable to our policy and follow the same training pipeline. Additionally, LOTUS utilizes the DINO-v2 (Oquab et al., 2024) foundation model instead of CLIP, as recommended by the ablation study in their paper.
  • Task-specific Adapters for Imitation Learning (TAIL) (Liu et al., 2024b) introduces separate LoRAs into the spatial encoder and temporal decoder for each new task. The fusion module, state encoder, and action head are also full fine-tuned and stored as task-specific adapters. TAIL requires an oracle task identifier to retrieve the corresponding adapters during inference. In order to align with other methods, we additionally provide the result of SeqFT (LoRA), which fine-tunes LoRA parameters sequentially.
  • Learning to Modulate (L2M) (Schmied et al., 2023) maintains a learnable modulation pool with keys and associated modulators. The embedded history of state tokens serves as the query vector for query-key matching, which finally steers the pretrained policy’s behavior. We follow the original paper to use a frozen fusion module and action head, and insert LoRA pools in the CLIP encoders and the temporal transformer backbone. We use a pool size of 30 and the similarity loss coefficient $\lambda=0.5$.
  • Incremental Learning of Retrievable Skills for Continual Imitation Learning (IsCiL) (Lee et al., 2024) establishes multifaceted prototypes through K-means clustering in the state space, where proper LoRA-based skills is retrieved upon current input state. We define the task-specific skill the same as in TAIL (Liu et al., 2024b), i.e., LoRA matrices for transformer, the fusion module, the state encoder, and the action head. We use CLIP encoded features as context embeddings and set the skill prototype $\mathcal{X}_{z}$ in $\mathcal{X}$ to be composed of 20 bases.
Table 1: Policy Configurations
Hyperparameter Value
@@
Pretrained model CLIP ViT-B/16
Image resolution 128*128
@@
Joint state dim. 7
Gripper state dim. 2
@@
Number of hidden layers 1
Hidden size 256
@@
Number of layers 6
Number of attn. heads 8
Embedding size 768
MLP hidden size 1024
Max sequence Length 10
Dropout 0.15
@@
Number of hidden layers 2
Hidden size 256
Number of Modes 5
Min. std 1e-4
Activation Softplus
Action dim. 7
@@
Number of hidden layers 2
Hidden Activation GELU
@@
Table 2: Training Hyperparameters
Hyperparameter Value
Optimizer
Training epochs 10
Batch size 32
Optimizer AdamW
Betas (0.9,0.999)
Learning rate 1e-4
Anneal strategy cos
Weight decay 0.1
Gradient clip 100
Image Augmentation
Brightness 0.3
Saturation 0.3
Contrast 0.3
Hue 0.3
Color jitter prob. 0.9
Rotation 15
Translation 0.1
Affine prob. 0.6
Evaluation
Epochs per eval. 2
Eval. episodes 20
Algorithm 1 Training Process of DMPEL
0: Pretrained policy $\pi_{\theta}$, router $\mathcal{R}$, low-rank expert library $\mathcal{L}=\varnothing$ for each layer $\mathcal{F}$, coefficient replay buffer $\mathcal{B}_{\text{CR}}=\varnothing$
1:for $k=1,2,\cdots,K$ do
2:  for each $\{\mathcal{F},\mathcal{L}\}$ do
3:   Initialize $\bm{A}_{k}$, $\bm{B}_{k}$, (and optionally $\bm{b}_{k}$)
4:   $\mathcal{L}\leftarrow\text{sg}(\mathcal{L})\cup\{\bm{A}_{k},\bm{B}_{k},\bm{b}_{k}\}$ # $\text{sg}(\cdot)$ means stop gradient, i.e., freezing previously learned low-rank experts
5:  end for
6:  for epochs $c=1,2\cdots,C$ do
7:   for each mini-batch data from dataset $\mathcal{D}^{k}$ do
8:    Compute the context embedding $\bm{r}$ using the frozen encoders in the pretrained policy $\pi_{\theta}$
9:    Compute the expert coefficient vector $\bm{c}$ with the router $\mathcal{R}$ using Eq. (5)
10:    for each $\{\mathcal{F},\mathcal{L}\}$ do
11:     Obtain the expert coefficient $\bm{c}^{\diamond}$ according to which sub-module $\diamond$ the layer belongs to
12:     Synthesize parameters $\tilde{\bm{W}},\tilde{\bm{b}}$ using Eq. (6)
13:    end for
14:    Compute behavior cloning loss $L$ using Eq. (1)
15:    Update router parameters $\mathcal{R}$ and low-rank expert parameters $\bm{A}_{k}$, $\bm{B}_{k}$, (and optionally $\bm{b}_{k}$) to minimize $L$
16:    if $c=C$ then $\mathcal{B}_{\text{CR}}\leftarrow\mathcal{B}_{\text{CR}}\cup\{(\bm{r},\bm{c})\}$ with a probability equals to the coefficient replay ratio
17:   end for
18:   if $c\in\mathcal{C}$ then Evaluate on task $\mathcal{T}^{k}$ to obtain success rate $s_{k,k,c}$ (Algorithm 2)
19:  end for
20:  for epochs $c=1,2\cdots,C$ do
21:   for each mini-batch data from the buffer $\mathcal{B}_{\text{CR}}$ do
22:    Compute expert coefficient replay loss $L_{\text{CR}}$ using Eq. (7)
23:    Update router parameters $\mathcal{R}$ to minimize $L_{\text{CR}}$
24:   end for
25:  end for
26:  for $j=1,2,\cdots,k$ do
27:   Evaluate on task $\mathcal{T}^{j}$ to obtain success rate $s_{k,j}$ (Algorithm 2)
28:  end for
29:end for
30: Compute metrics FWT, NBT, AUC using Eq. (2)-(4)
Algorithm 2 Inference Process of DMPEL
0: Pretrained policy $\pi_{\theta}$, router $\mathcal{R}$, each layer with pretrained weight and low-rank expert library $\{\mathcal{F},\mathcal{L}\}$, FIFO queues $F_{q}=\varnothing$, $R_{q}=\varnothing$ with maximum length $T$
1:for $t=1,2,\cdots,\mathcal{H}$ do
2:  Observe $o_{t}=(I^{1}_{t},\cdots,I^{N_{c}}_{t},p_{t})$ and language instruction $l$
3:  if $T_{\text{syn}}\mid t$ then
4:   Compute visual and textual embeddings $\tilde{f}^{v}_{t}$, $\tilde{f}^{l}_{t}$ using the frozen CLIP encoders $\mathcal{E}_{I}$, $\mathcal{E}_{L}$ ($\mathcal{F}$ only)
5:   Compute the context embedding $\bm{r}_{t}=[\tilde{f}^{v}_{t},\tilde{f}^{l}_{t},p_{t}]$
6:   $R_{q}\leftarrow R_{q}\cup\{\bm{r}_{t}\}$
7:   Compute the expert coefficient vector $\bm{c}_{t}=\text{top-k}(\mathcal{R}(\bm{r}_{t-T:t}),\delta)$
8:   for each $\{\mathcal{F},\mathcal{L}\}$ do
9:    Obtain the expert coefficient $\bm{c}^{\diamond}_{t}$ according to which sub-module $\diamond$ the layer belongs to
10:    Synthesize parameters $\tilde{\bm{W}},\tilde{\bm{b}}$ using Eq. (6)
11:   end for
12:  else
13:   Reuse previous visual and text embeddings $\tilde{f}^{v}_{t}\leftarrow\tilde{f}^{v}_{t-1}$, $\tilde{f}^{l}_{t}\leftarrow\tilde{f}^{l}_{t-1}$
14:   $R_{q}\leftarrow R_{q}\cup\{[\tilde{f}^{v}_{t}$, $\tilde{f}^{l}_{t},p_{t}]\}$
15:  end if
16:  Compute visual, textual, and state embeddings $f^{v}_{t}$, $f^{l}_{t}$, $f^{s}_{t}$ using encoders $\mathcal{E}_{I}$, $\mathcal{E}_{L}$, $\mathcal{E}_{P}$ ($\mathcal{F}$ and $\mathcal{L}$)
17:  Obtain $f_{t}\in\mathbb{R}^{\text{num of modalities}\times\text{embedding dim}}$ using FiLM to fuse language embeddings with image and state embeddings
18:  $F_{q}\leftarrow F_{q}\cup\{f_{t}\}$
19:  Compute latent embedding $g_{t}=\mathcal{D}_{T}(f_{t-T},\cdots,f_{t-1},f_{t})$
20:  Predict action $a_{t}=\mathcal{D}_{H}(g_{t})$
21:  Interact with the environment using action $a_{t}$ and obtain done flag $d_{t+1}$
22:  if $d_{t+1}\text{ is True}$ then break
23:end for
We utilize the implementations of SeqFT, ER, EWC, and PackNet provided in LIBERO (Liu et al., 2024a) and we directly use the official code of LOTUS (Wan et al., 2024). When training with ER and EWC, we use two GPUs, each with a batch size of 16, ensuring the total batch size remains consistent with other methods. Due to the unavailability of publicly released code for TAIL, we implemented it based on the description in the paper (Liu et al., 2024b). We integrated the code for L2M and IsCiL into our codebase, as the original implementations operate with different policies in different benchmarks.
For LoRA-based methods, we insert LoRA modules with rank $r=8$ into the 6-th to 11-th layers of the CLIP image and text encoders, and LoRA modules with $r=16$ into the temporal transformer. LoRA is applied to the query and value projection matrices. In TAIL and IsCiL, we learn task-specific fusion modules, state encoders, and action heads, as described in (Liu et al., 2024b). In contrast, these components are frozen in L2M following the original paper (Schmied et al., 2023).
DMPEL (Ours).
We adopt the same settings in the image encoder, text encoder, and temporal transformer as other LoRA-based methods, and use LoRA experts with $r=16$ for all linear layers in other sub-modules. Besides, in the Long suite, we tune both the low-rank weights $\bm{A},\bm{B}$ and biases $\bm{b}$, whereas in the Goal, Spatial, and Object suites, we only tune the low-rank weights $\bm{A},\bm{B}$. However, both settings lead to fewer learnable parameters than TAIL (Liu et al., 2024b), which learns these sub-modules separately for each task. We also employ dropout with probability $p=0.15$ on the coefficient for low-rank experts to avoid overfitting.
We provide a graphic overview of DMPEL in Figure 2 in the main text and the detailed pseudocode in Algorithm 1 (Training) and 2 (Inference). For further implementation specifics, please refer to the source code.
@@ A.2 Training Configuration
We generally follow the setup in the LIBERO benchmark for training and evaluation (Liu et al., 2024a). For all experiments, we use either Nvidia A100 or H800 GPUs. We employ Distributed Data Parallel (DDP) for parallel training across 4 GPUs during pretraining, while using a single GPU for lifelong learning. We employ 16-bit floating point precision (FP16) for training and inference to accelerate the process, with the exception of the router, which uses FP32. This selective precision technique is also discussed in (Fedus et al., 2022). To ensure reproducibility and statistical significance, we adopted three random seeds (100, 200, 300) and reported the mean and standard deviation of the performance over three seeds throughout the paper.
We deviate from the original LIBERO benchmark by training each task for only 10 epochs (evaluating every 2 epochs), instead of 50 epochs (evaluating every 10 epochs), as we observed convergence typically within 10 epochs. Consequently, metrics including FWT and AUC that are averaged over epochs generally appear lower than those reported in the original LIBERO paper and some follow-up work. We also explored various hyperparameter settings for DMPEL, such as top-k selection, replay ratio coefficients, and weight initialization; these ablations are detailed in Figure 5. Otherwise, we use the same training hyperparameters across all methods and benchmarks, which proved effective. Hyperparameter details are summarized in Table 2.
Table 3: Tasks Instructions of LIBERO Task Suites
Benchmark Suite Task Instructions
@@
LIBERO-Goal put the bowl on top of the cabinet
@@
LIBERO-Spatial pick up the black bowl in the top drawer of the wooden cabinet and place it on the plate
@@
LIBERO-Object pick up the ketchup and place it in the basket
@@
LIBERO-Long put the white mug on the left plate and put the yellow and white mug on the right plate
@@
A.3 Benchmark Details
We present the specific language instructions for the tasks in four lifelong learning LIBERO task suites in Table 3. Although some tasks have similar descriptions, they are not identical due to differences in environment configurations, such as varying spatial layouts, objects, or goal positions. Each task provides 50 successful expert demonstrations for imitation learning.
@@ Appendix B Additional Results
B.1 Ablation Studies
((a))
((b))
((c))
Figure 8: Ablation studies in the LIBERO-Goal suite. (a) Different task orders (b) Different rank sizes (c) Different number of demonstrations.
Figure 9: Ablation studies on router design in the LIBERO-Goal suite.
We further investigate the performance of DMPEL across various settings. First, we assess its robustness to different task orders, as shown in Figure 8(a). The default order provided by the LIBERO benchmark is 0123456789, the reverse order is 9876543210, and a randomly generated task order is 4687312095. Additionally, we analyze the impact of the rank size of each LoRA expert and the number of demonstrations for each new task, illustrated in Figures 8(b) and 8(c), respectively.
We utilize the Sigmoid function to constrain the router output instead of applying Softmax normalization. As noted in the related work, DMPEL is heavily inspired by research on model fusion and merging, such as SD-LoRA (Wu et al., 2025) and LoRAHub (Huang et al., 2024), which demonstrate that adjusting only the magnitude of LoRA can yield promising transfer results. The interval [0, 2] serves as a hyperparameter based on intuition to limit the magnitude, with 1 as the midpoint, representing the direct use of the previous LoRA module without magnitude alteration. We empirically compare the performance of Sigmoid and Softmax with various top-k strategies, as shown in Figure 9. Notably, router with Sigmoid generally outperforms the ones with Softmax, and the top-1 with Softmax variant exhibited the lowest performance. We hypothesize that using raw coefficients to modulate the magnitudes of LoRA experts provides greater flexibility, as these coefficients are independent for each expert. In contrast, Softmax normalization enforces that the sum of coefficients equals 1, meaning that increasing one coefficient necessarily decreases another. During expert coefficient replay, we apply the MSE loss between old and new coefficients in all cases for its simplicity. When transitioning from task $k$ to task $k+1$, we pad the coefficient vector with 0 to maintain alignment with the current library size and ensure that later introduced experts are excluded from previous tasks. We hypothesize that minimizing the discrepancy of logits may be a more effective approach for the top-1 Softmax variant due to its one-hot property. In this context, padding zeros to the coefficients becomes padding -$\infty$ to the logits. We leave further investigation of this for future work.
Beyond the experiment on the standard $K=10$ task sequence form the LIBERO benchmark, we also investigate the applicability of DMPEL to ultra-long sequences. To this end, we concatenate the LIBERO-Object, Goal, and Spatial benchmark to form a task sequence with $K=30$. As mentioned in Section 5.2 and illustrated in Figure 6(b), we prune underactivated low-rank experts every 10 tasks to control the growth of the library. The results shown in Figure 10 indicate that DMPEL consistently outperforms FFT with experience replay in success rate, while incurring significantly lower computational and storage costs (see Figure 1). However, we observe a gradual decline in performance as the number of tasks increases. Given the small size of the router, this is not surprising. When the number of distinct tasks continue to increase, the router may be exhausted. One potential solution is to progressively scale up the router as the task count increases. Since all expert coefficients are stored in the replay buffer, we should be able to recover the coefficient distribution in the enlarged router. We plan to explore this in future research.
Figure 10: Average success rate on encountered tasks in the ultra long task sequence concatenated with LIBERO-Object, Goal, and Spatial benchmarks.
B.2 Visualization Analysis
We plot some representative trajectories from different benchmark suites in Figure 11-13. In Figure 11, we observe that the temporal decoder reuses the previously learned low-rank experts throughout the entire trajectory, while the state encoder relies on a newly introduced expert, especially when grasping the bowl. In Figure 12, we can see the switching process between different experts as the robot approaches, grasps, transports, and places down the bowl. Both the vision and state encoder relies on the new expert. In Figure 13, the vision encoder reuses previously learned expert but the fusion sub-module learns a new expert to effectively fuse features for manipulation.
((a))
((b))
Figure 11: Visualization Analysis on task 9 from LIBERO-Goal task: put the bowl on the plate.
((a))
((b))
Figure 12: Visualization Analysis on task 7 from LIBERO-Spatial task: pick up the black bowl next to the cookie box and place it on the plate.
((a))
((b))
Figure 13: Visualization Analysis on task 5 from LIBERO-Object task: pick up the ketchup and place it in the basket.
B.3 Learning curves
We present the learning curves for all methods in four lifelong learning LIBERO task suites in Figure 14-17 . The $y$-axis indicates the success rate, while the $x$-axis depicts the agent’s training process over the course of lifelong learning. For example, the Task 1 subplot in the figure shows the agent’s performance on the first task when it learns ten tasks sequentially. Learning curves demonstrate that DMPEL achieves superior forward transfer with reduced forgetting compared to baselines.
@@ B.3 Learning curves
Appendix C The Usage of Large Language Models
We only use Large Language Models (LLMs) to assist and polish writing based on prompts such as "Please help me polish the paragraph and check for any grammatical mistakes."